# When Algorithms Meet Artists: Semantic Compression of Artists’ Concerns in the Public AI-Art Debate

### v3

@@ 1.1 The Paradoxical Position of Artists
Artists occupy a paradoxical position in this technological transition. They stand as both the primary authors of training data for these systems and the stakeholders most susceptible to the systems’ disruptive impacts. Generative models are fundamentally reshaping the production, circulation, valuation, and attribution of creative work, while also intensifying persistent disputes concerning consent, compensation, authorship, and the legitimacy of data practices Lovato and others [2024], AOI Coalition of Artists [2025], Creators Coalition on AI [2025]. The effects of these disputes have become increasingly tangible. In 2023, a landmark class-action lawsuit (Andersen v. Stability AI) brought artists’ grievances before a federal court, alleging that generative image models developed using copyrighted works without consent constitute mass infringement Andersen et al. v. Stability AI et al. [2023]. As Varvasovszky and Brugha Varvasovszky and Brugha [2000] emphasize, identifying and understanding primary stakeholders is essential for evaluating the social consequences of technological change. For artists, this transformation raises fundamental questions about creative ownership, labor, and artistic identity McCormack et al. [2019], Browne [2022].
@@ 1.2 Stakeholder Frames: The “Territory” of Artist Concerns
Relevant research shows that we cannot treat “artists” as a single, unified group in their interpretation of AI developments. Studies by Lovato et al. Lovato and others [2024] and Kawakami & Venkatagiri Kawakami and Venkatagiri [2024] reveal that artist outlooks are actually highly specific and often internally complex. Lovato et al. Lovato and others [2024] operationalize these concerns across five dimensions (Utility, Ownership, Compensation, Transparency, and Threat) comprising 34 distinct “frames,” understood here in Entman’s Entman [1993] sense as interpretive packages that select and emphasize particular aspects of a perceived reality. This granularity is essential: one artist may embrace AI tools while rejecting the economic models of companies that develop them; another might support transparency requirements but remain uncertain about ownership claims. Similarly, Kawakami and Venkatagiri Kawakami and Venkatagiri [2024] document the spectrum of impacts that generative AI has on creative practitioners, ranging from the benefits of workflow augmentation to the considerable employment risks that accompany its widespread implementation, while Jiang et al. Jiang and others [2023] identify distinct patterns of concerns across different artistic communities.
@@ 1.3 Public Discourse: The “Map” that Compresses the Territory
Zai et al. Zai and others [2025] examined AI debates across Western media, finding that economic and scientific actors possess disproportionate standing, systematically foregrounding narratives of “progress” and economic potential while overshadowing ethical and social frames. Similarly, Bøgh Bøgh [2025] found that news media tend to reduce the complex AI-art interface into binary oppositions—“AI as Threat” versus “AI as Tool”—stripping away the nuanced labor concerns held by practitioners. Banks and Li Banks and Li [2025] extend this analysis by mapping public debates about image-generative AI specifically, revealing how particular framings achieve dominance while others remain peripheral.
The structural mechanics of the contemporary public sphere exacerbates this phenomenon. Sichach Sichach [2025] argues that we have transitioned from traditional editorial gatekeeping to “algorithmic agenda-setting,” where recommender systems prioritize content that maximizes engagement rather than epistemic completeness. Pane Pane [2025] describes this as a “post-mediatized” sphere in which algorithmic curation establishes the epistemic boundaries of discourse, determining which voices are amplified and which remain invisible. Huang and Gadavanij Huang and Gadavanij [2025] demonstrate that in AI-related media coverage, not all social actors are represented equally, and limited coverage contributes to systematic underrepresentation of affected stakeholders. Together, the evidence suggests that the social and structural conditions exist for the systematic marginalization of artist voices—a hypothesis that has not been empirically tested in the AI-art domain using computational methods that can quantify the degree and nature of this marginalization.
@@ 1.4 Epistemic Injustice: Marginalization Without Explicit Exclusion
Recent scholarship on epistemic injustice in AI governance provides a useful theoretical lens. Baeyaert Baeyaert [2025] demonstrates that when knowledge produced by affected groups is systematically backgrounded relative to institutional or technical narratives, marginalization can occur even without explicit exclusion—through differential amplification, topic selection, and standing asymmetries rather than overt silencing. In the AI-art debate, where artists are directly affected by data practices and market restructuring, epistemic marginalization is a plausible outcome of the visibility regimes documented by Sichach Sichach [2025] and Pane Pane [2025]. The question is not whether artists are formally prohibited from speaking, but whether their concerns achieve adequate salience within the discursive ecosystem that shapes public understanding and policy attention.
@@ 1.5 Temporal Context: From Novelty to Crisis
The timing of this question matters because AI-art discourse has shifted sharply over the last decade. Early public discussion (2013–2020) disproportionately emphasized novelty, experimentation, and philosophical speculation about machine creativity, framed by milestones such as Google’s DeepDream Mordvintsev et al. [2015] and the emergence of Generative Adversarial Networks Goodfellow et al. [2014], which spawned discussions of algorithmic aesthetics and machine consciousness. This period generated substantial scholarly attention to questions of algorithmic aesthetics, machine consciousness, and the ontological status of AI-generated works Browne [2022], Mazzone and Elgammal [2019]. Mazzone and Elgammal Mazzone and Elgammal [2019] explored the creative potential of AI systems, while Browne Browne [2022] interrogated the very category of “AI artist.” The 2018 Christie’s auction of Portrait of Edmond de Belamy—a GAN-generated image—brought AI art into mainstream cultural discourse, intensifying debates about authorship and market legitimacy.
The more recent period (2021–2025), characterized by the mass adoption of text-to-image systems like DALL-E Ramesh et al. [2021] and latent diffusion models Rombach et al. [2022], has seen a pivot toward labor disruption, legal conflict, and institutional integration. Cheung Cheung [2024] shows that media environments frequently oscillate between “opportunity” and “crisis” narratives, an alternation that can simplify stakeholder realities into a small set of repeatable tropes. Saeedi and Taleghani Saeedi and Taleghani [2025] apply topic modeling to examine the discourses surrounding generative art diffusion, revealing how particular themes achieve prominence at different temporal moments. However, their analysis focuses on the internal structure of public discourse without systematically comparing it to empirically-derived stakeholder positions. Despite this shift, we lack systematic empirical evidence on the alignment between public narratives and stakeholder priorities: does public discourse reflect the distribution of artists’ concerns, or does it compress them into a narrow semantic region?
@@ 1.6 Research Gap: Semantic Compression as Measurable Marginalization
This gap is especially important for computational social science because it connects a normative claim—that artists are primary stakeholders whose voices should inform AI governance—to a measurable outcome: representation and salience within public discourse. This gap also connects to broader questions in behavioral science about how collective attention is allocated and how stakeholder knowledge is aggregated (or fails to aggregate) in public deliberation. Existing research often focuses on technological capabilities, philosophical questions of machine creativity, or legal analysis of copyright Saeedi and Taleghani [2025], Banks and Li [2025], Kawakami and Venkatagiri [2024], Mazzone and Elgammal [2019], Jiang and others [2023], but seldom measures how stakeholder concerns are geometrically distributed within the broader discursive ecosystem. As DiMaggio, Nag, and Blei DiMaggio et al. [2013] argue, topic modeling offers a powerful tool for exploiting affinities between computational text analysis and the sociological study of culture, enabling researchers to map meaning structures at scale while remaining attentive to the relational positions of different actors within those structures.
@@ 1.7 This Study: Mapping Discourse, Projecting Stakeholders
In this paper, we address this gap by mapping the semantic landscape of public discourse on AI and art (2013–2025) and projecting artist survey statements into that space to quantify alignment, separation, and salience. Our public corpus spans 131 documents and 891 processed text chunks across news, podcasts, panel discussions, legal filings, and research papers—yielding 22 distinct discourse topics. We operationalize stakeholder frames using the five dimensions identified by Lovato et al. Lovato and others [2024], comprising 34 distinct frames from 252 US-based practicing artists.
@@
    H2: Distributional divergence between artist frames and public discourse will persist after style controls, indicating semantic, not stylistic, marginalization.
@@ 2.1 Study Design Overview
We measured whether artist stakeholder concerns achieve proportional representation in public discourse about generative AI and art. Our approach constructs a semantic reference map from public discourse (2013–2025) and projects artist survey responses into that shared space to quantify representational alignment—an approach aligned with recent computational methods for comparing how different stakeholder groups articulate contested topics.
We operationalize artist concerns using five dimensions from prior survey research Lovato and others [2024]: Threat, Utility, Transparency, Ownership, and Compensation. These dimensions structure our extraction of artist probes, generation of style-matched comparison probes, and interpretation of theme-specific misalignment.
@@ 2.2.1 Public Discourse Corpus
We included English-language sources from news articles, blogs, podcasts, interviews, panel discussions, peer-reviewed papers, and legal filings. We excluded social-media-first platforms (Reddit, Twitter/X) to maintain comparability with search-ranked and editorialized sources Sichach [2025], Pane [2025]. Spoken sources were transcribed using OpenAI Whisper with manual review.
@@ 2.2.2 Artist Survey Corpus
Artist stakeholder language was derived from the survey dataset of Lovato et al. Lovato and others [2024], which captured practicing artists’ attitudes toward generative AI. We filtered to US-based respondents actively practicing as artists ($n=252$) to match the geographic scope of our public corpus.
@@ 2.2.3 Public Probes (Style Control)
Using GPT-5.1, we generated 250 synthetic Likert-style anchor statements spanning all five themes, then retrieved the nearest-neighbor sentences from the public corpus. A human-LLM validation filter retained only substantively on-theme sentences, yielding 379 public probes. This enables direct comparison: if artist-public divergence persists when both are expressed as short declaratives, the mismatch is semantic rather than stylistic.
@@ 2.4 Topic Interpretation
Each cluster was interpreted independently by five annotators: two human coders (first author and research assistant) and three LLMs (GPT-5.1 Pro, Gemini Pro 1.5, Claude Opus 4.5). This multi-annotator approach addresses subjectivity in topic labeling. A final judge (GPT-5.1 Pro) consolidated interpretations into final labels, keywords, and macro-thematic groupings (see Supplementary Table S2 for full topic inventory).
@@ 2.7 Use of Large Language Models
We used LLMs (GPT-5.1 Pro, Gemini Pro 1.5, Claude Opus 4.5) for anchor generation, validation filtering, and cluster interpretation. In all cases, human oversight was retained, and we employed multiple models to mitigate single-model biases. Artist survey data are available through Lovato et al. Lovato and others [2024].
@@ 3.1 Overview
We constructed a semantic map of public AI-art discourse from 131 documents (891 text chunks) spanning 2013–2025, including news articles, podcasts, panel discussions, academic papers, and legal filings. Clustering identified 22 distinct topics organized into five macro-thematic groups: Technical Genealogy (18.5%), Institutions and Markets (35.6%), Governance and Rights (26.0%), Philosophy of Creativity (7.5%), and Practice and Pedagogy (12.4%). The largest group—Institutions and Markets—reflects the prominence of curatorial, commercial, and spectacular narratives in public visibility. Full topic descriptions are provided in Supplementary Table S2.
@@ 3.2 Artist Concerns Concentrate in a Narrow Discursive Region (H1)
Figure 1: Artist concerns concentrate in a narrow region of public discourse. Distribution of 1,259 artist probes across 22 public discourse topics. Just 4 topics capture 95.4% of artist concerns, while 14 topics (62.4% of public discourse volume) contain no artist perspective whatsoever. The 34 distinct frames articulated by artists—representing substantively different positions on ownership, transparency, compensation, threat, and utility—collapse into this narrow semantic region.
@@
This concentration is not an artifact of limited artist diversity. Artists in our sample articulate 34 distinct frames representing substantively different positions: some embrace AI as a creative tool while others view it as an existential threat; some would donate their work freely while others demand ongoing royalties; some assert exclusive ownership rights while others reject property frameworks entirely. Yet this heterogeneity—the “territory” of stakeholder reality—collapses into a narrow region of the public discourse “map.”
@@ 3.4 Governance Concerns Are Most Marginalized (H3)
Not all artist concerns are equally underrepresented. We computed salience ratios—the concentration of artist probes in theme-relevant regions divided by the concentration of public discourse in those same regions—to quantify differential marginalization across concern dimensions (Figure 2).
Theme Salience Ratio (vs. Public) Salience Ratio (vs. Public Probes)
@@ 3.5 Summary
The 34 distinct frames articulated by artists—representing substantively different positions on threat, utility, transparency, ownership, and compensation—collapse into a narrow region of public meaning-space. Meanwhile, institutional, technical, and spectacular narratives occupy the broader discursive landscape. In the Discussion, we interpret this pattern as a form of epistemic marginalization with implications for AI governance.

### v4

@@ 1.1 The Paradoxical Position of Artists
Artists occupy a paradoxical position in this technological transition. They stand as both the primary authors of training data for these systems and the stakeholders most susceptible to the systems’ disruptive impacts. Generative models are fundamentally reshaping the production, circulation, valuation, and attribution of creative work, while also intensifying persistent disputes concerning consent, compensation, authorship, and the legitimacy of data practices (Lovato et al., 2024; AOI Coalition of Artists, 2025; Creators Coalition on AI, 2025). The effects of these disputes have become increasingly tangible. In 2023, a landmark class-action lawsuit (Andersen v. Stability AI) brought artists’ grievances before a federal court, alleging that generative image models developed using copyrighted works without consent constitute mass infringement (Andersen et al. v. Stability AI et al., 2023). As Varvasovszky and Brugha (2000) emphasize, identifying and understanding primary stakeholders is essential for evaluating the social consequences of technological change. For artists, this transformation raises fundamental questions about creative ownership, labor, and artistic identity (McCormack et al., 2019; Browne, 2022).
@@ 1.2 Stakeholder Frames: The “Territory” of Artist Concerns
Relevant research shows that we cannot treat “artists” as a single, unified group in their interpretation of AI developments. Studies by Lovato et al. (2024) and Kawakami and Venkatagiri (2024) reveal that artist outlooks are actually highly specific and often internally complex. Lovato et al. (2024) operationalize these concerns across five dimensions (Utility, Ownership, Compensation, Transparency, and Threat) comprising 34 distinct “frames,” understood here in Entman’s (1993) sense as interpretive packages that select and emphasize particular aspects of a perceived reality. This granularity is essential: one artist may embrace AI tools while rejecting the economic models of companies that develop them; another might support transparency requirements but remain uncertain about ownership claims. Similarly, Kawakami and Venkatagiri (2024) document the spectrum of impacts that generative AI has on creative practitioners, ranging from the benefits of workflow augmentation to the considerable employment risks that accompany its widespread implementation, while Jiang et al. (2023) identify distinct patterns of concerns across different artistic communities.
@@ 1.3 Public Discourse: The “Map” that Compresses the Territory
Zai et al. (2025) examined AI debates across Western media, finding that economic and scientific actors possess disproportionate standing, systematically foregrounding narratives of “progress” and economic potential while overshadowing ethical and social frames. Similarly, Bøgh (2025) found that news media tend to reduce the complex AI-art interface into binary oppositions—“AI as Threat” versus “AI as Tool”—stripping away the nuanced labor concerns held by practitioners. Banks and Li (2025) extend this analysis by mapping public debates about image-generative AI specifically, revealing how particular framings achieve dominance while others remain peripheral.
The structural mechanics of the contemporary public sphere exacerbates this phenomenon. Sichach (2025) argues that we have transitioned from traditional editorial gatekeeping to “algorithmic agenda-setting,” where recommender systems prioritize content that maximizes engagement rather than epistemic completeness. Pane (2025) describes this as a “post-mediatized” sphere in which algorithmic curation establishes the epistemic boundaries of discourse, determining which voices are amplified and which remain invisible. Huang and Gadavanij (2025) demonstrate that in AI-related media coverage, not all social actors are represented equally, and limited coverage contributes to systematic underrepresentation of affected stakeholders. Together, the evidence suggests that the social and structural conditions exist for the systematic marginalization of artist voices—a hypothesis that has not been empirically tested in the AI-art domain using computational methods that can quantify the degree and nature of this marginalization.
@@ 1.4 Epistemic Injustice: Marginalization Without Explicit Exclusion
Recent scholarship on epistemic injustice in AI governance provides a useful theoretical lens. Baeyaert (2025) demonstrates that when knowledge produced by affected groups is systematically backgrounded relative to institutional or technical narratives, marginalization can occur even without explicit exclusion—through differential amplification, topic selection, and standing asymmetries rather than overt silencing. In the AI-art debate, where artists are directly affected by data practices and market restructuring, epistemic marginalization is a plausible outcome of the visibility regimes documented by Sichach (2025) and Pane (2025). The question is not whether artists are formally prohibited from speaking, but whether their concerns achieve adequate salience within the discursive ecosystem that shapes public understanding and policy attention.
@@ 1.5 Temporal Context: From Novelty to Crisis
The timing of this question matters because AI-art discourse has shifted sharply over the last decade. Early public discussion (2013–2020) disproportionately emphasized novelty, experimentation, and philosophical speculation about machine creativity, framed by milestones such as Google’s DeepDream (Mordvintsev et al., 2015) and the emergence of Generative Adversarial Networks (Goodfellow et al., 2014), which spawned discussions of algorithmic aesthetics and machine consciousness. This period generated substantial scholarly attention to questions of algorithmic aesthetics, machine consciousness, and the ontological status of AI-generated works (Browne, 2022; Mazzone and Elgammal, 2019). Mazzone and Elgammal (2019) explored the creative potential of AI systems, while Browne (2022) interrogated the very category of “AI artist.” The 2018 Christie’s auction of Portrait of Edmond de Belamy—a GAN-generated image—brought AI art into mainstream cultural discourse, intensifying debates about authorship and market legitimacy.
The more recent period (2021–2025), characterized by the mass adoption of text-to-image systems like DALL-E (Ramesh et al., 2021) and latent diffusion models (Rombach et al., 2022), has seen a pivot toward labor disruption, legal conflict, and institutional integration. Cheung (2024) shows that media environments frequently oscillate between “opportunity” and “crisis” narratives, an alternation that can simplify stakeholder realities into a small set of repeatable tropes. Saeedi and Taleghani (2025) apply topic modeling to examine the discourses surrounding generative art diffusion, revealing how particular themes achieve prominence at different temporal moments. However, their analysis focuses on the internal structure of public discourse without systematically comparing it to empirically-derived stakeholder positions. Despite this shift, we lack systematic empirical evidence on the alignment between public narratives and stakeholder priorities: does public discourse reflect the distribution of artists’ concerns, or does it compress them into a narrow semantic region?
@@ 1.6 Research Gap: Semantic Compression as Measurable Marginalization
This gap is especially important for computational social science because it connects a normative claim—that artists are primary stakeholders whose voices should inform AI governance—to a measurable outcome: representation and salience within public discourse. This gap also connects to broader questions in behavioral science about how collective attention is allocated and how stakeholder knowledge is aggregated (or fails to aggregate) in public deliberation. Existing research often focuses on technological capabilities, philosophical questions of machine creativity, or legal analysis of copyright (Saeedi and Taleghani, 2025; Banks and Li, 2025; Kawakami and Venkatagiri, 2024; Mazzone and Elgammal, 2019; Jiang et al., 2023), but seldom measures how stakeholder concerns are geometrically distributed within the broader discursive ecosystem. As DiMaggio et al. (2013) argue, topic modeling offers a powerful tool for exploiting affinities between computational text analysis and the sociological study of culture, enabling researchers to map meaning structures at scale while remaining attentive to the relational positions of different actors within those structures.
@@ 1.7 This Study: Mapping Discourse, Projecting Stakeholders
In this paper, we address this gap by mapping the semantic landscape of public discourse on AI and art (2013–2025) and projecting artist survey statements into that space to quantify alignment, separation, and salience. Our public corpus spans 131 documents and 891 processed text chunks across news, podcasts, panel discussions, legal filings, and research papers—yielding 22 distinct discourse topics. We operationalize stakeholder frames using the five dimensions identified by Lovato et al. (2024), comprising 34 distinct frames from 252 US-based practicing artists.
@@
    H2: Distributional divergence between artist frames and public discourse will persist after style controls, indicating semantic marginalization.
@@ 2.1 Study Design Overview
We measured whether artist stakeholder concerns achieve proportional representation in public discourse about generative AI and art. Our approach constructs a semantic reference map from public discourse (2013–2025) and projects artist survey responses into that shared space to quantify representational alignment—an approach aligned with recent computational methods for comparing how different stakeholder groups articulate contested topics (Elmholdt et al., 2025; Matsui and Ferrara, 2024).
We operationalize artist concerns using five dimensions from prior survey research (Lovato et al., 2024): Threat, Utility, Transparency, Ownership, and Compensation. These dimensions structure our extraction of artist probes, generation of style-matched comparison probes, and interpretation of theme-specific misalignment.
@@ 2.2.1 Public Discourse Corpus
We included English-language sources from news articles, blogs, podcasts, interviews, panel discussions, peer-reviewed papers, and legal filings. We excluded social-media-first platforms (Reddit, Twitter/X) to maintain comparability with search-ranked and editorialized sources (Sichach, 2025; Pane, 2025). Spoken sources were transcribed using OpenAI Whisper with manual review.
@@ 2.2.2 Artist Survey Corpus
Artist stakeholder language was derived from the survey dataset of Lovato et al. (2024), which captured practicing artists’ attitudes toward generative AI. We filtered to US-based respondents actively practicing as artists ($n=252$) to match the geographic scope of our public corpus.
@@ 2.2.3 Public Probes (Style Control)
Using GPT-5.1, we generated 250 synthetic Likert-style anchor statements following a factorial design: 5 themes $\times$ 5 agreement levels (strongly disagree to strongly agree) $\times$ 10 discourse-style variants (e.g., blog opinion, news editorial, artist interview, legal brief; see Supplementary Table S4 for design matrix and representative examples; full set available in our code repository). We then retrieved the nearest-neighbor sentences from the public corpus for each anchor. A human-LLM validation filter (Ziems et al., 2024) retained only substantively on-theme sentences, yielding 379 public probes. This enables direct comparison: if artist-public divergence persists when both are expressed as short declaratives, the mismatch is semantic rather than stylistic.
@@ 2.4 Topic Interpretation
Each cluster was interpreted independently by five annotators: two human coders (first author and research assistant) and three LLMs (GPT-5.1 Pro, Gemini Pro 1.5, Claude Opus 4.5). This multi-annotator approach addresses subjectivity in topic labeling (Ziems et al., 2024). A final judge (GPT-5.1 Pro) consolidated interpretations into final labels, keywords, and macro-thematic groupings (see Supplementary Table S2 for full topic inventory).
@@ 2.7 Use of Large Language Models
We used LLMs (GPT-5.1 Pro, Gemini Pro 1.5, Claude Opus 4.5) for anchor generation, validation filtering, and cluster interpretation. In all cases, human oversight was retained, and we employed multiple models to mitigate single-model biases (Ziems et al., 2024). Artist survey data are available through Lovato et al. (2024).
@@ 3.1 Overview
We constructed a semantic map of public AI-art discourse from 131 documents (891 text chunks) spanning 2013–2025, including news articles, podcasts, panel discussions, academic papers, and legal filings. Clustering identified 22 distinct topics organized into five macro-thematic groups (Table 1). The largest group—Institutions and Markets (35.6%)—reflects the prominence of curatorial, commercial, and spectacular narratives in public visibility. Full topic descriptions are provided in Supplementary Table S2.
Table 1: Macro-thematic groupings of the 22 public discourse topics, with share of public corpus and number of constituent topics.
Macro-theme Share of corpus Topics
Institutions & Markets 35.6% 7
Governance & Rights 26.0% 5
Technical Genealogy 18.5% 5
Practice & Pedagogy 12.4% 3
Philosophy of Creativity 7.5% 2
@@ 3.2 Artist Concerns Concentrate in a Narrow Discursive Region (H1)
Figure 1: Artist concerns concentrate in a narrow region of public discourse. Distribution of 1,259 artist probes across 22 public discourse topics. Just 4 topics capture 95.4% of artist concerns, while 14 topics (62.4% of public discourse volume) contain no artist perspective whatsoever.
@@
This concentration is not an artifact of limited artist diversity. Artists in our sample articulate 34 distinct frames representing substantively different positions: some embrace AI as a creative tool while others view it as an existential threat; some would donate their work freely while others demand ongoing royalties; some assert exclusive ownership rights while others reject property frameworks entirely. Yet this heterogeneity—the “territory” of stakeholder reality—collapses into a narrow region of the public discourse “map” (Figure 2).
Figure 2: Key mismatches between public discourse and artist concerns across the macro-thematic groups. Public discourse is dominated by Institutions & Markets (35.6%) and Technical Genealogy (18.5%), while artist concerns concentrate heavily in Governance & Rights (57.5%).
@@ 3.4 Governance Concerns Are Most Marginalized (H3)
Not all artist concerns are equally underrepresented. We computed salience ratios—the concentration of artist probes in theme-relevant regions divided by the concentration of public discourse in those same regions—to quantify differential marginalization across concern dimensions.
Theme SR (vs. Public) SR (vs. Public Probes)
@@ 3.5 Summary
The 34 distinct frames articulated by artists—representing substantively different positions on threat, utility, transparency, ownership, and compensation—collapse into a narrow region of public meaning-space. In the Discussion, we interpret this pattern as a form of epistemic marginalization with implications for AI governance.